OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Enter your answer as an ordered pair. y = 2x - 3 y = -2x + 5

OpenStudy (anonymous):

They intersect at (2,1)

OpenStudy (anonymous):

so is that right?

OpenStudy (anonymous):

equate both equation to find out where they intersect: 2x-3=-2x+5 4x=8 x=2 no pug in x into any one of the equations to find y: y=2(2)-3= 4-3=1

OpenStudy (anonymous):

i got (1.5,0)

OpenStudy (amistre64):

y = 2x - 3 ; *-1 y = -2x + 5 -y = -2x +3 y = -2x +5 ----------- 0 = -4x +8 ; solve for x -8 = -4x x = 2

OpenStudy (anonymous):

apex check you work, amistre and I are getting the same values

OpenStudy (anonymous):

(2,1) ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about this one y = -x + 5 y = x - 3

OpenStudy (anonymous):

x-3=-x+5 2x=8 x=4 y=4-3 y=1

OpenStudy (anonymous):

y = -x + 2 y = -2x + 6

OpenStudy (anonymous):

-x+2= -2x+6 x+2=6 x=4 y=-(4)+2= -2 (4,-2)

OpenStudy (anonymous):

u = -b + 21 u = -2b + 30

OpenStudy (anonymous):

(9, 12)

OpenStudy (anonymous):

you are my new buddy help me on this one

OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. y = 9x + 4 x = 6

OpenStudy (anonymous):

hello